# In-Context Learning Is Provably Bayesian Inference: A Generalization Theory for Meta-Learning

### v1

@@ Appendix H Proofs of the Technical Lemmas
Step 2 (Localization at worst–path sequential radius). Let $h_{\theta}:=M_{\theta}-M_{\mathrm{Bayes}}$. Fix a $\mathcal{Z}$}–valued predictable tree $Z=(Z_{k})_{k=1}^{p}$ of depth $p$ that is decoupled tangent to the prompt process, in the sense that for each depth $k$ and each past $\xi_{1:k-1}\in\{\pm 1\}^{k-1}$, the conditional distribution of $Z_{k}(\xi_{1:k-1})$ equals the conditional distribution of $P^{k}$ given $D^{k-1}$ (Peña & Giné, 1999; Rakhlin et al., 2015); namely $Z_{k}(\xi_{1:k-1})\mid D^{k-1}\stackrel{{\scriptstyle d}}{{=}}P^{k}\mid D^{k-1}$. Conditioning on a realization $Z=z$, we refer to such $z$ as a data-containing tangent tree. For any realization $z$ of $Z$, define the worst–path sequential $\ell_{2}$ radius on $z$ by
@@
We decompose $\widetilde{\Lambda}_{1}(\theta)=A_{1}(\theta)+B_{1}(\theta)$. From the definition of ${\mathcal{H}}(r)$, $\mathbb{E}[\sup_{\theta\in{\mathcal{H}}(r)}A_{1}^{2}(\theta)]^{1/2}\leq r^{2}$. We then analyze $B_{1}^{2}(\theta)=\{\frac{2}{p}\sum_{k=1}^{p}(y_{1,k+1}-M_{\mathrm{Bayes}}(P_{1}^{k}))(M_{\mathrm{Bayes}}(P_{1}^{k})-M_{\theta}(P_{1}^{k}))\}^{2}$. Note that $|M_{\mathrm{Bayes}}(P_{1}^{k})-M_{\theta}(P_{1}^{k})|\leq B_{f}+B_{M}$ and $B_{1}$ is constructed by a martingale difference sequence with filtration ${\mathcal{G}}^{\prime}_{k}$. Since $\mathbb{E}[X^{2}]=2\int_{0}^{\infty}t\Pr(|X|>t)\mathrm{d}t\leq 2\int_{t_{0}}^{\infty}t\Pr(|X|>t)\mathrm{d}t+2\int_{0}^{t_{0}}t\mathrm{d}t$, evaluation of the tail probability from Lemma 8 in Rakhlin et al. (2015) yields
@@
Step 5 (Sequential Dudley bound). The sequential Dudley integral bound (Block et al., 2021, Corollary 10) gives, for an absolute constant $C>0$,
@@
where $N^{\prime}$ denotes the fractional covering number (Block et al., 2021). Note that since every $h\in{\mathcal{H}}(r)$ satisfies $\|h\|_{\mathrm{seq},2;z}\leq r$, the diameter under the path $\ell_{2}$ metric is at most $2r$, so the upper limit can be replaced by $2r$, from Lemma 7 in Block et al. (2021),
@@
By the standard symmetrization, we have $\mathbb{E}\sup_{\theta\in\Theta(r)}\big|(\widehat{\mathcal{R}}-\mathcal{R})(\theta)-(\widehat{\mathcal{R}}-\mathcal{R})(\theta^{\star})\big|\lesssim\frac{1}{\sqrt{N}}\{\mathfrak{R}^{\mathrm{seq}}_{p}\big(\mathcal{H}(r)\big)+\sqrt{\,r+R_{\mathrm{BG}}(M_{\theta^{\star}})\,}\}$. Then, from Lemma 4 and Corollary 10 in Block et al. (2021), there exists a constant $c>0$ (range rescaling absorbed into $c$) such that $\mathbb{E}\sup_{\theta\in\Theta(r)}\!\big|(\widehat{\mathcal{R}}-\mathcal{R})(\theta)-(\widehat{\mathcal{R}}-\mathcal{R})(\theta^{\star})\big|\lesssim\sqrt{\frac{r+R_{\mathrm{BG}}(M_{\theta^{\star}})}{N}}\big(1+\sqrt{\log\frac{c_{1}}{r+R_{\mathrm{BG}}(M_{\theta^{\star}})}}+\sqrt{\frac{m}{p}}\;\sqrt{\log\frac{c_{2}\sqrt{m}}{r+R_{\mathrm{BG}}(M_{\theta^{\star}})}}\big).$
@@
Appendix I LLM Usage Disclosure

### v2

@@ Appendix I Proofs of the Technical Lemmas
Step 2 (Localization at worst–path sequential radius). Let $h_{\theta}:=M_{\theta}-M_{\mathrm{Bayes}}$. Fix a $\mathcal{Z}$}–valued predictable tree $Z=(Z_{k})_{k=1}^{p}$ of depth $p$ that is decoupled tangent to the prompt process, in the sense that for each depth $k$ and each past $\xi_{1:k-1}\in\{\pm 1\}^{k-1}$, the conditional distribution of $Z_{k}(\xi_{1:k-1})$ equals the conditional distribution of $P^{k}$ given $D^{k-1}$ (pena1999decoupling; Rakhlin2015); namely $Z_{k}(\xi_{1:k-1})\mid D^{k-1}\stackrel{{\scriptstyle d}}{{=}}P^{k}\mid D^{k-1}$. Conditioning on a realization $Z=z$, we refer to such $z$ as a data-containing tangent tree. For any realization $z$ of $Z$, define the worst–path sequential $\ell_{2}$ radius on $z$ by
@@
We decompose $\widetilde{\Lambda}_{1}(\theta)=A_{1}(\theta)+B_{1}(\theta)$. From the definition of ${\mathcal{H}}(r)$, $\mathbb{E}[\sup_{\theta\in{\mathcal{H}}(r)}A_{1}^{2}(\theta)]^{1/2}\leq r^{2}$. We then analyze $B_{1}^{2}(\theta)=\{\frac{2}{p}\sum_{k=1}^{p}(y_{1,k+1}-M_{\mathrm{Bayes}}(P_{1}^{k}))(M_{\mathrm{Bayes}}(P_{1}^{k})-M_{\theta}(P_{1}^{k}))\}^{2}$. Note that $|M_{\mathrm{Bayes}}(P_{1}^{k})-M_{\theta}(P_{1}^{k})|\leq B_{f}+B_{M}$ and $B_{1}$ is constructed by a martingale difference sequence with filtration ${\mathcal{G}}^{\prime}_{k}$. Since $\mathbb{E}[X^{2}]=2\int_{0}^{\infty}t\Pr(|X|>t)\mathrm{d}t\leq 2\int_{t_{0}}^{\infty}t\Pr(|X|>t)\mathrm{d}t+2\int_{0}^{t_{0}}t\mathrm{d}t$, evaluation of the tail probability from Lemma 8 in Rakhlin2015 yields
@@
Step 5 (Sequential Dudley bound). The sequential Dudley integral bound (BlockDaganRakhlin2021, Corollary 10) gives, for an absolute constant $C>0$,
@@
where $N^{\prime}$ denotes the fractional covering number (BlockDaganRakhlin2021). Note that since every $h\in{\mathcal{H}}(r)$ satisfies $\|h\|_{\mathrm{seq},2;z}\leq r$, the diameter under the path $\ell_{2}$ metric is at most $2r$, so the upper limit can be replaced by $2r$, from Lemma 7 in BlockDaganRakhlin2021,
@@
By the standard symmetrization, we have $\mathbb{E}\sup_{\theta\in\Theta(r)}\big|(\widehat{\mathcal{R}}-\mathcal{R})(\theta)-(\widehat{\mathcal{R}}-\mathcal{R})(\theta^{\star})\big|\lesssim\frac{1}{\sqrt{N}}\{\mathfrak{R}^{\mathrm{seq}}_{p}\big(\mathcal{H}(r)\big)+\sqrt{\,r+R_{\mathrm{BG}}(M_{\theta^{\star}})\,}\}$. Then, from Lemma 4 and Corollary 10 in BlockDaganRakhlin2021, there exists a constant $c>0$ (range rescaling absorbed into $c$) such that $\mathbb{E}\sup_{\theta\in\Theta(r)}\!\big|(\widehat{\mathcal{R}}-\mathcal{R})(\theta)-(\widehat{\mathcal{R}}-\mathcal{R})(\theta^{\star})\big|\lesssim\sqrt{\frac{r+R_{\mathrm{BG}}(M_{\theta^{\star}})}{N}}\big(1+\sqrt{\log\frac{c_{1}}{r+R_{\mathrm{BG}}(M_{\theta^{\star}})}}+\sqrt{\frac{m}{p}}\;\sqrt{\log\frac{c_{2}\sqrt{m}}{r+R_{\mathrm{BG}}(M_{\theta^{\star}})}}\big).$
@@
Appendix J LLM Usage Disclosure